A tour group split into two groups when waiting in line for food at a fast food counter. The first group bought 8 slices of pizza and 7 soft drinks for $34.29. The second group bought 5 slices of pizza and 7 soft drinks for $24.87. How much does one slice of pizza cost?
Answer by mananth(16949)

cost/unit
pizza x
soft drink y
Total 750
8 x + 7 y = 34.29 .............1
5 x + 7 y = 24.87 .............2
Eliminate y
multiply (1)by -1
Multiply (2) by 1
-8 x -7 y = -34.29
5 x + 7 y = 24.87
Add the two equations
-3 x = -9.42
/ -3
x = 3.14
plug value of x in (1)
8 x + 7 y = 34.29
25.12 + 7 y = 34.29
7 y = 34.29 -25.12
7 y = 9.17
y = 1.31
pizza $ 3.14
soft drink $ 1.31
m.ananth@hotmail.ca